A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 108.6-cm and a standard deviation of 2.5-cm. For shipment, 24 steel rods are bundled together.

Find the probability that the average length of a randomly selected bundle of steel rods is less than 108.5-cm.
P(M < 108.5-cm) =


Enter your answer as a number accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

\sigma\bar x = \sigma / \sqrt n = 2.5 / \sqrt 24

P(M < 108.5) = P((M - \mu M ) / \sigma M < (108.5 - 108.6) / 2.5 / \sqrt 24)

= P(z < -0.196)

P(M < 108.5) = 0.4223